# Coupled Twiss Parameters Estimation from TbT Data

### v2

@@ I Introduction
Methods utilizing TbT data are also widely used for the determination of uncoupled optical functions [9, 10, 11]. Although not as accurate as the methods based on the ORM [12], these methods provide fast optics inference. In these methods, optical functions can be estimated using the amplitude or phase of coherent oscillations derived from the harmonic analysis of TbT signals. Typically, the ideal linear model is designed to be uncoupled, and the corresponding optical functions are estimated without coupling. However, estimating coupled linear Twiss parameters is also important, especially when the initial accelerator model includes coupling by design or when the lattice contains sources of undesired coupling that need identification. For example, the SKIF light source [13], currently under construction, will feature a special operation mode that uses coupling to increase the beam lifetime [14]. Similar to methods based on the closed orbit response, methods based on TbT signals are applicable for examining the coupled optical model.
In this paper, we consider several methods for estimating the linear coupled Twiss parameters of transverse motion. To achieve this, we construct the normalization matrix at each BPM location, transforming the original phase space coordinates into normalized ones. In this normalized space, the motion is uncoupled, and the phase space trajectories are circular. BPMs provide information only about the beam centroid coordinates, but not the momenta. We explore several methods for reconstructing transverse momenta in a linear approximation [15, 16]. With both coordinates and momenta, it is feasible to approximate the one-turn matrix at each BPM [17]. For each fitted matrix, we perform a symplectification procedure [18] and determine the corresponding parameters of the coupled optical model [19]. It is noteworthy that Twiss parameters can be computed by fitting not just the one-turn matrix but also its power, potentially enhancing the signal-to-noise ratio [17].
The paper introduces two more methods for estimating coupled Twiss parameters. These methods involve approximating normalization matrices using spectra of complex normalized coordinates and linear coupled invariants. They use a special form of the normalization matrix that does not contain singularities in the limit of zero coupling. All the methods discussed in this article rely on the computation of reconstructed transverse momenta using two or more BPMs. This makes them dependent on the BPM calibration accuracy.
@@ II Uncoupled Twiss Parameters
The modern BPM systems enable the measurement of transverse TbT coordinates with high accuracy. If coherent oscillations around the closed orbit are excited, for example, by a pulsed electromagnetic field, then the signal from BPM $i$ can be used to estimate the frequency $\nu$, amplitude $a_{i}$, and phase $\varphi_{i}$ of the oscillations. In the linear uncoupled case, oscillations around the closed orbit can be described as follows
@@
where $I$ is the linear action, and $\beta_{i}$ is the beta function at BPM $i$. Measured TbT amplitude values can be used to estimate $\beta_{i}$ value, provided the value of the action $I$ is known. The action itself can be estimated using either model values or values measured by a different method of the beta function $\bar{\beta}_{i}$, and measured amplitude values as an average value over all BPMs
@@
The beta functions at each BPM can be estimated using approximated action $\hat{I}$ as follows
$\beta_{i}=\frac{a_{i}^{2}}{\hat{I}}$ (3)
@@
Amplitudes allow the estimation of only beta functions, which are of primary interest but do not provide a complete description of the motion. The Courant-Snyder parameters $\alpha$ and $\beta$ can be used for the full parameterization of linear uncoupled motion. The transition matrix between locations $i$ and $j$ can be represented using these parameters as
@@
Figure 1: Different cases of phase advance computation from $i$ to $j$. Phase advance is considered negative if propagation from $i$ to $j$ implies moving back in time. If BPMs are located on different turns, the tune should be adjusted before computing the phase advance.
Figure 1 illustrates the computation of phase advance from location $i$ to $j$, given the total number of BPMs $n$, where $i$ and $j$ can be any integers. Given a location index $i\in\mathbb{Z}$, its corresponding monitor index is defined as $\bar{i}=i\mod n$, i.e. index on the first turn. With the measured phase advance $\varphi_{\bar{i}}$ at this index and a frequency $\nu$, the phase is given by
$\varphi_{i}=\mod\left(\varphi_{\bar{i}}+2\pi\nu k,2\pi\right),\quad k=\lfloor\frac{i-\bar{i}}{n}\rfloor$ (5)
Using the above expression, the phase advance from location $i$ to $j$ is expressed as
@@
The normalization matrix $N$ is symplectic and depends on the Twiss parameters as follows
@@
For this particular form of the normalization matrix, given monitors $i$ and $j$, and the transport matrix $M_{i,j}$ between them, the following identity holds
@@
The identity above provides an equation that relates Twiss parameters at location $i$. To determine the parameters $\alpha_{i}$ and $\beta_{i}$, at least two additional monitors, apart from the probed one, are required. This requirement forms the basis of the well-known triplet method, as defined in [9]. Model transport matrices can be utilized to estimate measured Twiss parameters. It should be noted that the measurable values of $\alpha$ and $\beta$ are conveniently expressed through the measured phase advance, model values $\alpha^{m}$ and $\beta^{m}$, and model phase advance as follows
@@
where $\alpha^{m}i$ and $\beta^{m}i$ are the model Twiss parameters at the probed monitor $i$, $\mu^{m}{ij}$ and $\mu^{m}{ik}$ are the model phase advances from $i$ to $j$ and $k$, respectively, while $\mu_{ij}$ and $\mu_{ik}$ are the corresponding measured phase advances.
@@ III Coupled Twiss Parameters
Different parameterizations, as referenced in [27, 28, 29], are used to transform linear coupled motion into its normal form. While most parameterizations primarily focus on the transverse planes, this paper utilizes the parameterization proposed by Wolski [19]. The main objective of any parameterization is to construct a symplectic transformation. This transformation converts the original system, which includes coupling, into a normalized uncoupled system, where the phase trajectories form circles
@@
where $R$ is the rotation matrix corresponding to the normalized system, $M$ is the one-turn matrix of the original system, $N$ is the normalization matrix. $X_{n}$ and $X_{n+1}$ are the normalized coordinates at turn $n$ and $n+1$, respectively. The matrix $N$ must be symplectic and comply with the chosen gauge. For instance, the condition $n_{i,i+1}\equiv 0,~{}i=2k+1$ and $n_{i,i}>0$ can be applied to establish a direct relationship between the $n_{i,i}$ elements and the beam sizes.
In Wolski’s paper, the normalization matrix $N$ is derived from the eigenvectors of the one-turn matrix $M$
@@
where the eigenvectors $e_{k}$ are normalized such that $e_{k}^{H}Se_{k^{\prime}}=i\delta_{k,k^{\prime}}$ and $S$ is the standard symplectic form
@@
It is important to note that the one-turn matrix $M$ and the matrix $\Sigma S$ (where $\Sigma$ is a transformation matrix for beam second moments) share the same eigenvectors. The transformation of beam second moments is governed by the equation
$\Sigma_{j}=M_{i,j}\Sigma_{i}M_{i,h}^{T}$ (13)
@@
The normalizing matrix $N$ is constructed from the eigenvectors of the one-turn matrix $M$ as follows
$N=EQ=\left(e_{-I}~{}|~{}e_{I}~{}|~{}e_{-II}~{}|~{}e_{II}\right)\\ \begin{pmatrix}\hat{Q}&0\\ 0&\hat{Q}\end{pmatrix}$ (14)
and the $\hat{Q}$ matrix is constructed from eigenvectors of rotation matrix and is given by
@@
Here, the eigenvectors $e_{-I}$ and $e_{I}$ correspond to the first degree of freedom, while $e_{-II}$ and $e_{II}$ correspond to the second degree of freedom. For each degree of freedom $k$, Twiss matrices $B^{k}$ are introduced
@@
where matrices $T^{k}$ perform the projection, for example, for two degrees of freedom
@@
The Twiss parameter matrices $B^{k}$ are symmetric by construction and are transformed similarly to the matrix of the second moments $\Sigma$
@@
In Wolski’s article, a method for estimating coupled Twiss parameters from phase data is presented. Equations similar to 8 in Section II are also valid in the coupled case with a modified normalization matrix $N_{i}$ and a with full $4\times 4$ transport matrix $M_{i,j}$ between monitors. Thus, is it possible to estimate coupled Twiss parameters using a larger number of BPMs. However, this method has several limitations. One of them is the requirement to use a large area around the probed monitor. Additionally, model transport matrices typically do not contain coupling by design, which results in the inability to compute coupling-specific Twiss parameters.
This latter problem becomes evident when examining Equation II in Section II. The value of $\beta$ is proportional to the model value $\beta^{m}_{i}$. In coupled case, one would expect a similar equation for coupled beta function. However, if its model value is zero, calculating the measured value becomes infeasible since the measured coupled beta function is directly proportional to it. Similarly, using model matrices without coupling doesn’t allow to estimate Twiss parameters specific to coupling.
@@ IV Parametric Normalization Matrix
The simplicity condition reduces the number of free elements of $4\times 4$ matrix to $10$, and $2$ elements, $n_{1,2}$ and $n_{3,4}$, are fixed to zero by the standard gauge. Additionally, the conditions $n_{1,1}>0$ and $n_{3,3}>0$ are assumed. The normalization matrix then has the following form
@@
where the parameters fixed by symplectic condition are denoted by the symbol $X$ and the elements $n_{1,1}>0$, $n_{1,2}\equiv 0$, $n_{1,3}$, $n_{1,4}$, $n_{2,1}$, $n_{3,1}$, $n_{3,3}>0$, $n_{3,4}\equiv 0$, $n_{4,1}$, and $n_{4,3}$, other elements are determined by solving algebraic equations obtained from symplectic condition $N^{T}SN=S$ with $S$ given by Equation 12 in Section III. This parametric normalization matrix can be used to relate linear invariants in measured and normalized coordinates. Different forms of coupled Twiss parameters can be computed from normalization matrix. The free elements for a particular normalization matrix used in this paper are given by
@@ V Transverse Momenta Reconstruction
BPMs do not provide direct measurements of transverse momenta. However, values of transverse momenta are necessary for estimating the coupled Twiss parameters with the methods described in this paper. Given the coordinates at two monitors and the model transformation between them, one can compute the transverse momenta at these monitors. The reconstruction procedure based on Taylor series representation of the transformation can be performed in the general nonlinear case [15, 16, 30]. In the linear case, transverse momenta can be expressed by explicit formulas. For instance, using the coordinates from monitors $i$ and $j$, the momenta at monitor $i$ are equal to
@@
where $m^{(i,j)}_{q,p}$ are elements of the transport matrix from monitor $i$ to $j$.
Given a pair of BPMs along with a model or a measured transport matrix between them, it is possible to reconstruct the values of transverse momenta. To improve the accuracy of transverse momenta estimation, multiple BPM pairs can be utilized. Combining results from different monitor pairs helps obtaining a more accurate and reliable estimate of momenta. Results from different pairs can be combined with weights, for example, using TbT noise estimate from signals without excitation or directly from oscillations [33]. Alternatively, the least squares (LS) method can also be used for transverse momenta estimation. With LS, both the coordinates and momenta are determined using the following system
@@
where the coordinates and momenta indexed by $i$ are fitted using several monitors within the selected range $j\in{(i-k,i+k)},~{}k\in\mathbb{N}^{+}$.
It should be noted that the reconstructed momenta depend on the calibrations of BPMs. The linear part of BPM calibrations can be described by a calibration matrix
$\begin{pmatrix}\bar{x}\\ \bar{y}\end{pmatrix}=G\begin{pmatrix}x\\ y\end{pmatrix}=\begin{pmatrix}g_{x,x}&g_{x,y}\\ g_{y,x}&g_{y,y}\end{pmatrix}\begin{pmatrix}x\\ y\end{pmatrix}$ (23)
where $g_{x,x}$ and $g_{y,y}$ are the calibration scales, $g_{x,y}$ and $g_{y,x}$ are calibration cross-talks, $x$ and $y$ correspond to actual coordinates, and $\bar{x}$ and $\bar{y}$ correspond to measured coordinates at a BPM. Following the approach outlined in [34], calibrations can be included as additional linear transformations at the BPM’s entrance and exit. The calibration matrix that transforms both coordinates and momenta, which is also symplectic, is given by
$\bar{G}=\begin{pmatrix}G&0\\ 0&\left(G^{-1}\right)^{T}\end{pmatrix}$ (24)
Consequently, the observed transport matrix between BPMs $i$ and $j$ is given by
$\bar{M}_{i,j}=\bar{G}_{j}M_{i,j}\bar{G}_{i}$ (25)
This illustrates how calibration errors can modify the observed twiss parameters and, hence, the importance of calibration independent methods based on phase measurements. Since methods based on reconstructed momenta depend on BPM calibrations, the estimated Twiss parameters correspond to the observed ones.
To account for uncertainties in transverse momenta estimation and to improve statistical reliability, both the bootstrap procedure and the Monte Carlo (MC) method can be used. The bootstrap procedure enables the assessment of random coordinate measurement errors by generating multiple samples (with replacement) from the available data. This approach provides a distribution of possible momentum values and an estimate of its mean value and corresponding standard errors. The MC method accounts for systematic errors in transport matrices by simulating random variations in accelerator elements and BPM parameters. Both methods, bootstrap and MC, are powerful tools for accounting for various sources of uncertainty in estimating transverse momenta. In the case of VEPP-4M, inclusion of systematic errors is challenging because the magnets do not have individual power sources. Therefore, only the bootstrap method is applied to momenta or coupled Twiss estimation procedures.

### v3

# Coupled Twiss Parameters Estimation from Turn-by-Turn Data

@@ I Introduction
Methods utilizing TbT data are also widely used for the determination of uncoupled optical functions [9, 10, 11]. Although not always as accurate as the methods based on the ORM, like in the ESRF case [12], these methods provide fast optics inference. In these methods, optical functions can be estimated using the amplitude or phase of coherent oscillations derived from the harmonic analysis of TbT signals. Typically, the ideal linear model is designed to be uncoupled, and the corresponding optical functions are estimated without coupling. However, estimating coupled linear Twiss parameters is also important, especially when the initial accelerator model includes coupling by design or when the lattice contains sources of undesired coupling that need identification. For example, in [13] effects of undesired coupling are discussed for the LHC. The SKIF light source [14], currently under construction, will feature a special operation mode that uses coupling to increase the beam lifetime [15]. Similar to methods based on the closed orbit response, methods based on TbT signals are applicable for examining the coupled optical model [16, 17, 18, 19].
In this paper, we consider several methods for estimating the linear coupled Twiss parameters of transverse motion. To achieve this, we construct the normalization matrix at each BPM location, transforming the original phase space coordinates into normalized ones. In this normalized space, the motion is uncoupled, and the phase space trajectories are circular. BPMs provide information only about the beam centroid coordinates, but not the momenta. We explore several methods for reconstructing transverse momenta in a linear approximation [20, 21, 22]. With both coordinates and momenta, it is feasible to approximate the one-turn matrix at each BPM [18]. For each fitted matrix, we perform a symplectification procedure [23] and determine the corresponding parameters of the coupled optical model [16]. It is noteworthy that Twiss parameters can be computed by fitting not just the one-turn matrix but also its power, potentially enhancing the signal-to-noise ratio [18]. Further examples of linear coupled motion parameterizations are discussed in references [24, 25, 26, 27].
In addition to the method based on the one-turn matrix, this paper introduces two more methods for estimating coupled Twiss parameters. These methods involve approximating normalization matrices using spectra of complex normalized coordinates and linear coupled invariants. They use a special form of the normalization matrix that does not contain singularities in the limit of zero coupling. All the methods discussed in this article rely on the computation of reconstructed transverse momenta using two or more BPMs. This makes them dependent on the BPM calibration accuracy.
@@ II Uncoupled Twiss Parameters
The modern BPM systems enable the measurement of transverse TbT coordinates with high accuracy. If coherent oscillations around the closed orbit are excited, for example, by a pulsed electromagnetic field, then the signal from BPM $i$ can be used to estimate the frequency $\nu$, amplitude $a_{i}$, and phase $\varphi_{i}$ of the oscillations. In the linear uncoupled case, oscillations around the closed orbit can be described as follows:
@@
where $I$ is the linear action, and $\beta_{i}$ is the beta function at BPM $i$. Measured TbT amplitude values can be used to estimate $\beta_{i}$ value, provided the value of the action $I$ is known. The action itself can be estimated using either model values or values measured by a different method of the beta function $\bar{\beta}_{i}$, and measured amplitude values as an average value over all BPMs:
@@
The beta functions at each BPM can be estimated using approximated action $\hat{I}$ as follows:
$\beta_{i}=\frac{a_{i}^{2}}{2\hat{I}}$ (3)
@@
Amplitudes allow the estimation of only beta functions, which are of primary interest but do not provide a complete description of the motion. The Courant-Snyder parameters $\alpha$ and $\beta$ can be used for the full parameterization of linear uncoupled motion. The transition matrix between locations $i$ and $j$ can be represented using these parameters as:
@@
Figure 1: (Color) Different cases of phase advance computation from $i$ to $j$. Phase advance is considered negative if propagation from $i$ to $j$ implies moving back in time. If BPMs are located on different turns, the tune should be adjusted before computing the phase advance.
Figure 1 illustrates the computation of phase advance from location $i$ to $j$, given the total number of BPMs $n$, where $i$ and $j$ can be any integers. Given a location index $i\in\mathbb{Z}$, its corresponding monitor index is defined as $\bar{i}=i\mod n$, i.e. index on the first turn. With the measured phase $\varphi_{\bar{i}}$ at this index and a frequency $\nu$, the phase is given by:
$\varphi_{i}=\mod\left(\varphi_{\bar{i}}+2\pi\nu k,2\pi\right),\quad k=\frac{i-\bar{i}}{n}$ (5)
Using the above expression, the phase advance from location $i$ to $j$ is expressed as:
@@
The normalization matrix $N$ is symplectic and depends on the Twiss parameters as follows:
@@
For this particular form of the normalization matrix (Equation 7), given monitors $i$ and $j$, and the transport matrix $M_{i,j}$ between them, the following identity holds:
@@
The identity above provides an equation that relates Twiss parameters at location $i$. To determine parameters $\alpha_{i}$ and $\beta_{i}$ from Equation 8, at least two additional monitors, apart from the probed one, are required. This requirement forms the basis of the well-known triplet method, as defined in [9]. Model transport matrices can be utilized to estimate measured Twiss parameters. It should be noted that the measurable values of $\alpha$ and $\beta$ are conveniently expressed through the measured phase advance, model values $\alpha^{m}$ and $\beta^{m}$, and model phase advance as follows:
@@
where $\alpha^{m}_{i}$ and $\beta^{m}_{i}$ are the model Twiss parameters at the probed monitor $i$, $\mu^{m}_{ij}$ and $\mu^{m}_{ik}$ are the model phase advances from $i$ to $j$ and $k$, respectively, while $\mu_{ij}$ and $\mu_{ik}$ are the corresponding measured phase advances.
@@ III Coupled Twiss Parameters
Different parameterizations, as referenced in [24, 25, 26, 27], are used to transform linear coupled motion into its normal form. While most parameterizations primarily focus on the transverse planes, this paper utilizes the parameterization proposed by Wolski [16]. The main objective of any parameterization is to construct a symplectic transformation. This transformation converts the original system, which includes coupling, into a normalized uncoupled system, where the phase trajectories form circles:
@@
where $R$ is the rotation matrix corresponding to the normalized system, $M$ is the one-turn matrix of the original system, $N$ is the normalization matrix. $X_{m}$ and $X_{m+1}$ are the normalized coordinates at turn $n$ and $n+1$, respectively. The matrix $N$ must be symplectic and comply with the chosen gauge. For instance, the condition $n_{i,i+1}\equiv 0,~{}i=2k+1$ and $n_{i,i}>0$ can be applied to establish a direct relationship between the $n_{i,i}$ elements and the beam sizes.
In Wolski’s paper, the normalization matrix $N$ is derived from the eigenvectors of the one-turn matrix $M$:
@@
where the eigenvectors $e_{k}$ are normalized such that $e_{k}^{H}Se_{k^{\prime}}=i\delta_{k,k^{\prime}}$ where $H$ denotes the Hermitian transpose and symplectic matrix $S$ is given by:
@@
It is important to note that the one-turn matrix $M$ and the matrix $\Sigma S$ (where $\Sigma$ is a transformation matrix for beam second moments) share the same eigenvectors. The transformation of beam second moments is governed by the equation:
$\Sigma_{j}=M_{i,j}\Sigma_{i}M_{i,j}^{T}$ (13)
@@
The normalizing matrix $N$ is constructed from the eigenvectors of the one-turn matrix $M$ as follows:
$N=EQ=\begin{pmatrix}e_{-I}&e_{I}&e_{-II}&e_{II}\end{pmatrix}\\ \begin{pmatrix}\hat{Q}&0\\ 0&\hat{Q}\end{pmatrix}$ (14)
and the $\hat{Q}$ matrix is constructed from eigenvectors of rotation matrix and is given by:
@@
Here, the eigenvectors $e_{-I}$ and $e_{I}$ correspond to the first degree of freedom, while $e_{-II}$ and $e_{II}$ correspond to the second degree of freedom. For each degree of freedom $k$, Twiss matrices $B^{k}$ are introduced:
@@
where matrices $T^{k}$ perform the projection, for example, for two degrees of freedom:
@@
The Twiss parameter matrices $B^{k}$ are symmetric by construction and are transformed similarly to the matrix of the second moments $\Sigma$:
@@
In Wolski’s article, a method for estimating coupled Twiss parameters from phase data is presented. Equations similar to 8 in Section II are also valid in the coupled case with a modified normalization matrix $N_{i}$ and with full $4\times 4$ transport matrix $M_{i,j}$ between monitors. Thus, is it possible to estimate coupled Twiss parameters using a larger number of BPMs. However, this method has several limitations. One of them is the requirement to use a large area around the probed monitor. Additionally, model transport matrices typically do not contain coupling by design, which results in the inability to compute coupling-specific Twiss parameters.
@@ IV Parametric Normalization Matrix
The symplecticity condition reduces the number of free elements of $4\times 4$ matrix to $10$, and $2$ elements, $n_{1,2}$ and $n_{3,4}$, are fixed to zero by the standard gauge. Additionally, the conditions $n_{1,1}>0$ and $n_{3,3}>0$ are assumed. The normalization matrix then has the following form:
@@
where the parameters fixed by symplectic condition are denoted by the symbol $X$ and the elements $n_{1,1}>0$, $n_{1,2}\equiv 0$, $n_{1,3}$, $n_{1,4}$, $n_{2,1}$, $n_{3,1}$, $n_{3,3}>0$, $n_{3,4}\equiv 0$, $n_{4,1}$, and $n_{4,3}$, other elements are determined by solving algebraic equations obtained from symplectic condition $N^{T}SN=S$ with $S$ given by Equation 12 in Section III. This parametric normalization matrix can be used to relate linear invariants in measured and normalized coordinates. Different forms of coupled Twiss parameters can be computed from normalization matrix. The free elements for a particular normalization matrix used in this paper are given by:
@@ V Transverse Momenta Reconstruction
BPMs do not provide direct measurements of transverse momenta. However, values of transverse momenta are necessary for estimating the coupled Twiss parameters with the methods described in this paper. Given the coordinates at two monitors and the model transformation between them, one can compute the transverse momenta at these monitors. The reconstruction procedure based on Taylor series representation of the transformation can be performed in the general nonlinear case [20, 21, 35]. In the linear case, transverse momenta can be expressed by explicit formulas. For instance, using the coordinates from monitors $i$ and $j$, the momenta at monitor $i$ are equal to:
@@
where $m^{(i,j)}_{q,p}$ are elements of the transport matrix from monitor $i$ to $j$. In the above, the mean values are assumed to be subtracted from the measured coordinates, prior to momenta computation, since Equations V and V are valid around the closed orbit.
Given a pair of BPMs along with a model or a measured transport matrix between them, it is possible to reconstruct the values of transverse momenta. To improve the accuracy of transverse momenta estimation, multiple BPM pairs can be utilized. Combining results from different monitor pairs helps obtaining a more accurate and reliable estimate of momenta. Results from different pairs can be combined with weights, for example, using TbT noise estimate from signals without excitation or directly from oscillations [38]. Alternatively, the least squares (LS) method can also be used for transverse momenta estimation. With LS, both the coordinates and momenta are determined using the following system:
@@
where the coordinates $(q_{x},q_{y})$ and momenta $(p_{x},p_{y})$ indexed by $i$ are fitted using several monitors within the selected range $j\in{(i-k,i+k)},~{}k\in\mathbb{N}^{+}$.
It should be noted that the reconstructed momenta depend on the calibrations of BPMs. The linear part of BPM calibrations can be described by a calibration matrix:
$\begin{pmatrix}\bar{q}_{x}\\ \bar{q}_{y}\end{pmatrix}=G\begin{pmatrix}q_{x}\\ q_{y}\end{pmatrix}=\begin{pmatrix}g_{x,x}&g_{x,y}\\ g_{y,x}&g_{y,y}\end{pmatrix}\begin{pmatrix}q_{x}\\ q_{y}\end{pmatrix}$ (23)
where $g_{x,x}$ and $g_{y,y}$ are the calibration scales, $g_{x,y}$ and $g_{y,x}$ are calibration cross-talks, $q_{x}$ and $q_{y}$ correspond to actual coordinates, and $\bar{q}_{x}$ and $\bar{q}_{y}$ correspond to measured coordinates at a BPM. Following the approach outlined in [39], calibrations can be included as additional linear transformations at the BPM’s entrance and exit. The symplectic calibration matrix that transforms both coordinates and momenta is given by:
$\bar{G}=\begin{pmatrix}g_{x,x}&g_{x,y}&0&0\\ 0&0&g_{y,y}/\det G&-g_{y,x}/\det G\\ g_{y,x}&g_{y,y}&0&0\\ 0&0&-g_{x,y}/\det G&g_{x,x}/\det G\end{pmatrix}$ (24)
Consequently, the observed transport matrix between BPMs $i$ and $j$ is given by:
$\bar{M}_{i,j}=\bar{G}_{j}M_{i,j}\bar{G}_{i}^{-1}$ (25)
This illustrates how calibration errors can modify the observed twiss parameters and, hence, the importance of calibration independent methods based on phase measurements. Since methods based on reconstructed momenta depend on BPM calibrations, the estimated Twiss parameters correspond to the parameters observed at the BPM frame. Transformation between BPM frame parameters and beam frame parameters can be performed using Equation 4, or 18.
To account for uncertainties in transverse momenta estimation and to improve statistical reliability, both the bootstrap procedure and the Monte Carlo (MC) method can be used. The bootstrap procedure enables the assessment of random coordinate measurement errors by generating multiple samples (with replacement) from the available data. This approach provides a distribution of possible momentum values and an estimate of its mean value and corresponding standard errors. The MC method accounts for systematic errors in transport matrices by simulating random variations in accelerator elements and BPM parameters. Both methods, bootstrap and MC, are powerful tools for accounting for various sources of uncertainty in estimating transverse momenta. In the case of VEPP-4M experimental data, only the bootstrap method is applied to momenta or coupled Twiss estimation procedures.